OpenStudy (anonymous):

Solve: -9(-3t + 8) = -13t + 48

OpenStudy (ganpat):

+27t - 72 = -13t + 48 27t + 13t = 48 + 72 40t = 120 t = 30...

OpenStudy (anonymous):

T can't = 30? You mean 3?

OpenStudy (ganpat):

No, i meant 30.. its explained above...

OpenStudy (anonymous):

Yeah, well that isn't a choice to choose from. I got 1, 3, -1, and 5

OpenStudy (ganpat):

oh sorry.................. its 3.. m extremely sorry...